OpenStudy (loser66):

Write gcd (210, 45) as a linear combination of 210 and 45 then use it to solve Diophantine 210x +45y = 60 Please, help

OpenStudy (trantom):

What does Diophantine mean?

OpenStudy (loser66):

My question is on second part. gcd (210, 45) =15 210 = 45(4) + 30 (a) 45= 30 + 15 (b) 30 = 15(2) +0 from (b) 15 = 45- 30 from (a) 30 = 210 +45(-4) hence 15 = 210 (-1) +45 (5) multiple both sides by 4, we get 60 = 210 (-4) +45 (20) hence \(x_0 = -4\\x = -4-3t ; t\in [0,14]\) \(y_0 = 20\\y = 20+14t; t\in[0,14]\)

OpenStudy (trantom):

I don't understand this yet sorry ganeshie can probably help :/

OpenStudy (loser66):

But if I use (a) to solve, I have 30 = 210 +45(-4) then multiple both sides by 2 to get 60 = 210 (2) +45(-8) I have different answer with \(x_0 = -2 \\y_0=-8\) it confused me a lot.

OpenStudy (loser66):

sorry \(x_0 =2\)

ganeshie8 (ganeshie8):

Both are correct

OpenStudy (loser66):

What?? explain me, please

ganeshie8 (ganeshie8):

\((2, -8)\) and \((-4, 20)\) are both particular solutions to the diophantine equation \( 210x +45y = 60\) You just need "any one" particular solution to write out all the solutions.

ganeshie8 (ganeshie8):

for example, \((-1, 6)\) is another particular solution

OpenStudy (loser66):

What is the logic?

ganeshie8 (ganeshie8):

all the solutions = (any one particular solution) + (null solutions)

OpenStudy (loser66):

null solution is fixed, right? how particular one changes without changing the whole?

ganeshie8 (ganeshie8):

Let me ask you a question, whats the vector equation of the line that passes through \((-4,20)\) and has a direction vector \((-3,14)\) ?

OpenStudy (loser66):

You do scare me, hehehe... I think it is a vector parallel to vector (-3,14) through (-4,20)

ganeshie8 (ganeshie8):

write out the equation, im sure you will see the connection once you write out the equaiton

ganeshie8 (ganeshie8):

whats the vector equation of the line that passes through \((-4,20)\) and has a direction vector \((-3,14)\) ? we can express the line as \[\dbinom{x}{y} = \dbinom{-4}{20} + t\dbinom{-3}{14}\] yes ?

OpenStudy (loser66):

it is y = -14/3x -4/3

ganeshie8 (ganeshie8):

I have asked for equation in "vector" form

OpenStudy (loser66):

ok,

OpenStudy (loser66):

yes, I got this part

ganeshie8 (ganeshie8):

look at the vector equation that i wrote is that the only way to express the equation of line in vector form ?

ganeshie8 (ganeshie8):

How about below equation ? \[\dbinom{x}{y} = \dbinom{2}{-8} + t\dbinom{-3}{14}\] what line does this represent ? does it represent a different line ?

OpenStudy (loser66):

Yes, they are same.

ganeshie8 (ganeshie8):

how do you know they are same ?

OpenStudy (loser66):

just starting point is different, the first one starts at (-4,20), the second one at (2,-8) and both are on the same line.

ganeshie8 (ganeshie8):

Exactly! so do you agree that we can use "any one point" on the line to express the line in vector form ?

OpenStudy (loser66):

Yes, I do. :)

ganeshie8 (ganeshie8):

good, so do you see how your solution for diophantine equation and the vector form of a line are similar ?

OpenStudy (loser66):

Yes, I got it. Thank you so much. You are.... amazingly patient.

ganeshie8 (ganeshie8):

In short : 1) (-4, 20) is a `particular solution` vector 2) all the `null solutions` are given by : t(-3, 14) 3) therefore `all the solutions` are given by (-4, 20) + t(-3, 14)

OpenStudy (loser66):

My midterm is upcoming. I need practice more. Do you know some practice pages online I can use to work with?

ganeshie8 (ganeshie8):

which textbook are you using ?

OpenStudy (loser66):

Elementary number theory by Kenneth H. Rosen

ganeshie8 (ganeshie8):

I have david m burton if you want more practice problems on linear diophantine equations, i can take a screenshot of exercise page and send you..

OpenStudy (loser66):

Thank you so much. For this time, I focus on : 1) The fundamental theorem 2)Linear Diophantine 3)Congruences 4) Wilson's theorem , Fermat's theorem and Euler's theorem.

OpenStudy (loser66):

one more: Chinese theorem also

ganeshie8 (ganeshie8):

Here are some fun problems. https://s3.amazonaws.com/upload.screenshot.co/962c284bf3 Solving these on your own gives solid understanding of diophantine equations...

OpenStudy (loser66):

Yes, I will.

ganeshie8 (ganeshie8):

Few more https://s3.amazonaws.com/upload.screenshot.co/c46dd4303d

OpenStudy (loser66):

I will try all. Thanks a lot.